Status Report

NOAA SATOPS Morning Report: Friday, August 13, 2010

By SpaceRef Editor
August 13, 2010
Filed under , , ,

Customers or users who want to call to ask about problems or report outages can use the following contacts:

– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 225

GOES
On DOY 224, GOES-13 operations were moved to WBU/Goddard at 1837 UTC and
GOES-11 to Fairbanks at 1810 UTC as a precaution due to flooding at Wallops. Wallops never lost power, but there was concern that water might rise enough to cause problems. The water has receded and a permanent fix is expected to be completed today. Planned fail back of operations to Wallops is DOY 225 1534 UTC for GOES-13 and 1550 UTC for GOES-11.

There were multiple GOES-12 framebreaks due to thunderstorms at Wallops.
GOES-11 images not collected due to eclipse: 0815, 0830, 0845, 0900 & 0930 UTC.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety

Questions or comments to Ron.Mahmot@noaa.gov

SpaceRef staff editor.